# How does non-metricity affect particle creation and evaporation in bumblebee gravity?

### v1

@@ Abstract
In this work, we analyze the impact of non–metricity on particle creation and the evaporation process of black holes within the framework of bumblebee gravity. In general lines, we compare black holes in the metric formalism 14 and the metric–affine approach filho2023vacuum . Initially, we focus on bosonic particle modes to investigate Hawking radiation. Using the Klein–Gordon equation, we compute the Bogoliubov coefficients and derive the Hawking temperature. Subsequently, we examine Hawking radiation as a tunneling process, resolving divergent integrals through the residue method. The analysis is then extended to fermionic particle modes, also within the tunneling framework. Particle creation densities are calculated for both bosonic and fermionic cases. Additionally, greybody bounds are estimated for bosonic and fermionic particles. Finally, we explore the evaporation process, considering the final state of the black holes. In a general panorama, non–metricity in bumblebee gravity raises particle density for bosons while reducing it for fermions, increases greybody factors (for both bosons and fermions), amplifies the emission rate, and accelerates the evaporation process.
@@ I Introduction
This study investigates how non–metricity influences particle creation and black hole evaporation in the context of bumblebee gravity. In order to address the question posed in the title of this manuscript, we specifically compare black holes described within the metric formalism 14 to those in the metric–affine framework filho2023vacuum . The analysis begins with bosonic particle modes, focusing on Hawking radiation. Scalar field is used to compute Bogoliubov coefficients and determine the Hawking temperature. Hawking radiation is also examined as a tunneling phenomenon, where divergent integrals are resolved using the residue method. The study then extends to fermionic particle modes, applying the tunneling approach as well to calculate particle creation densities. For bosons and fermions, greybody bounds are also evaluated. Finally, the evaporation process is analyzed, including the black hole’s final state. In general lines, non–metricity in bumblebee gravity increases particle density for bosons while decreasing it for fermions, raises greybody factors for both, intensifies the emission rate, and reduces the evaporation timescale.
@@ III.3.1 Scalar perturbations
Figure 4: The effective potential $\mathcal{V}^{\,\text{s}}_{{\text{metric}}}$ is shown for different values of $\ell$. Also, the Schwarzschild case is compared in this analysis.
@@ IV.3.2 Vector perturbations
Finally, in Fig. 21, the greybody factors for the bosonic case are compared. The following hierarchy is established: $\text{Schw}^{(\text{v})}>T_{b_{\text{met--aff}}}^{\text{v}}>T_{b_{\text{metric}}}^{\text{v}}>\text{Schw}^{(\text{s})}>T_{b_{\text{met--aff}}}^{\text{s}}>T_{b_{\text{metric}}}^{\text{s}}$.
@@
Figure 21: The comparison of the greybody factors for the bosonic case when $X=\ell=0.1$. The Schwarzschild case (for scalar and vectorial perturbations) is also present for the sake of comparison.

### v2

@@ Abstract
In this work, we analyze the impact of non–metricity on particle creation and the evaporation process of black holes within the framework of bumblebee gravity. In general lines, we compare black holes in the metric formalism 14 and the metric–affine approach filho2023vacuum . Initially, we focus on bosonic particle modes to investigate Hawking radiation. Using the Klein–Gordon equation, we compute the Bogoliubov coefficients and derive the Hawking temperature. Subsequently, we examine Hawking radiation as a tunneling process, resolving divergent integrals through the residue method. The analysis is then extended to fermionic particle modes, also within the tunneling framework. Particle creation densities are calculated for both bosonic and fermionic cases. Additionally, greybody bounds are estimated for bosonic and fermionic particles. Furthermore, we explore the evaporation process, considering the final state of the black holes and we also investigate the correlation between the greybody factors and the quasinormal modes. Finally, constraints on the Lorentz–violating parameters $\ell$ (for the metric case) and $X$ (for the metric–affine case) are established using recent astrophysical data on black hole lifetimes. In a general panorama, non–metricity (except for the tensor perturbations) in bumblebee gravity raises particle density for bosons while reducing it for fermions, increases greybody factors (for both bosons and fermions), amplifies the emission rate, and accelerates the evaporation process.
@@ I Introduction
The metric–affine formulation, also referred to as the Palatini approach, departs from the standard metric approach by treating the metric and affine connection as distinct geometric entities, each with its own dynamical role. This separation has led to a wide array of theoretical implications, as shown in several comprehensive studies Ghil1 ; Ghil2 ; gonzalo1 ; gonzalo2 ; gonzalo3 ; gonzalo4 . Nevertheless, investigations involving Lorentz symmetry breaking (LSB) remain limited within this framework. Recent contributions have started to bridge this gap, particularly by embedding spontaneous LSB mechanisms into the metric–affine setting through bumblebee gravity models Paulo2 ; Lambiase:2023zeo ; Jha:2023vhn ; Filho:2023etf ; Paulo3 ; AraujoFilho:2024ykw ; Paulo4 ; Filho:2022yrk .
This study investigates how non–metricity influences particle creation and black hole evaporation in the context of bumblebee gravity. In order to address the question posed in the title of this manuscript, we specifically compare black holes described within the metric formalism 14 to those in the metric–affine framework filho2023vacuum . The analysis begins with bosonic particle modes, focusing on Hawking radiation. Scalar field is used to compute Bogoliubov coefficients and determine the Hawking temperature. Hawking radiation is also examined as a tunneling phenomenon, where divergent integrals are resolved using the residue method. The study then extends to fermionic particle modes, applying the tunneling approach as well to calculate particle creation densities. For bosons and fermions, greybody bounds are also evaluated. Furthermore, the study also explores the evaporation process, emphasizing the end stage of black hole evolution. In addition, it examines how greybody factors are related to quasinormal modes. Constraints on the Lorentz–violating $\ell$ in the metric framework and $X$ in the metric–affine formulation—are derived by confronting the model with current astrophysical observations of black hole lifetime. In general lines, non–metricity (except for the tensor perturbations) in bumblebee gravity increases particle density for bosons while decreasing it for fermions, raises greybody factors for both, intensifies the emission rate, and reduces the evaporation timescale.
@@ III.3.1 Scalar perturbations
Figure 4: The effective potential $\mathcal{V}^{\,\text{s}}_{{\text{metric}}}$ is shown for different values of $\ell$ for $l=1$. Also, the Schwarzschild case is compared in this analysis.
@@ III.3.2 Vector perturbations
III.3.3 Tensor perturbations
The formulation of the master equations was carried out without anchoring the analysis to any specific foundational theory, relying solely on the presumed applicability of the Klein–Gordon and Maxwell dynamics. Yet, in gravitational settings where the matter sector is not minimally linked to the spacetime geometry described by $g_{\mu\nu}$, the usual conservation properties associated with these fields can break down.
Axial (odd–parity) gravitational perturbations demand variations not only in the gravitational field equations but also in the stress-energy content. When no fundamental theory is explicitly defined, a different strategy becomes necessary. Here, the analysis proceeds by interpreting the background through the lens of the Einstein equation, with modifications introduced via an effective stress–energy contribution. Comparable procedures have been employed across various scenarios in previous works AraujoFilho:2025rwr ; ashtekar2018quantum ; ashtekar2018quantum2 ; asasas2 ; baruah2025quasinormal . From a phenomenological standpoint, the stress–energy distribution corresponding to the black hole can be effectively described using an anisotropic fluid representation
$T_{\mu\nu}=\left(\rho+p_{2}\right)u_{\mu}u_{\nu}+\left(p_{1}-p_{2}\right)x_{\mu}x_{\nu}+p_{2}g_{\mu\nu}.$ (74)
In this setup, $\rho$ denotes the energy density measured in the rest frame of the fluid. The vector $u^{\mu}$ refers to the fluid’s timelike four–velocity, while $x^{\mu}$ is a unit spacelike vector orthogonal to both $u^{\mu}$ and the angular basis directions. Eq. (74) features $p_{1}$ and $p_{2}$, which characterize the pressures along the radial and tangential directions, respectively. Moreover, the vectors $u^{\mu}$ and $x^{\mu}$ obey the following normalization and orthogonality conditions:
$u_{\mu}u^{\mu}=-1\,,\qquad x_{\mu}x^{\mu}=1\,.$ (75)
The metric $g_{\mu\nu}$ serves as the operator for index manipulation throughout this analysis. In the rest frame of the fluid, the temporal and radial components of the basis vectors simplify to $u^{\mu}=(u^{t},0,0,0)$ for the four–velocity and $x^{\mu}=(0,x^{r},0,0)$ for the associated spacelike direction. Taking Eq. (75) into account, this configuration leads to the relation:
$u_{t}^{2}=g_{tt}(r)u_{t}u^{t}=-g_{tt}(r)\,,\qquad x_{r}^{2}=g_{rr}(r)x_{r}x^{r}=g_{rr}(r)\,.$ (76)
At the background level, the components of the stress–energy tensor take the following form:
$T_{tt}$ $=-g_{tt}(r)\rho\,,\qquad T_{t}^{t}=-\rho\,,$ (77)
$T_{rr}$ $=g_{rr}(r)p_{1}\,,\qquad T_{r}^{r}=p_{1}\,,$ (78)
$T_{\theta}^{\theta}=T_{\varphi}^{\varphi}=p_{2}\,.$ (79)
Notice that the radial profiles of $\rho$, $p_{1}$, and $p_{2}$ are determined by the structure of the spacetime and emerge from a direct computation of the Einstein tensor components.
To study the greybody facotrs in a static, spherically symmetric black hole background, a perturbation is introduced that alters the geometry into a time-dependent, axisymmetric form. This deformation leads to the following expression for the perturbed metric chen2019gravitational :
$\mathrm{d}s^{2}=$ $-e^{2\nu}\left(\mathrm{d}x^{0}\right)^{2}+e^{2\psi}\left(\mathrm{d}x^{1}-\sigma\mathrm{d}x^{0}-q_{2}\mathrm{d}x^{2}-q_{3}\mathrm{d}x^{3}\right)^{2}$
$+e^{2\mu_{2}}\left(\mathrm{d}x^{2}\right)^{2}+e^{2\mu_{3}}\left(\mathrm{d}x^{3}\right)^{2}\,.$ (80)
Here, it is important to highlight that the functions $\nu$, $\psi$, $\mu_{2}$, $\mu_{3}$, $\sigma$, $q_{2}$, and $q_{3}$ are taken to depend explicitly on the coordinates $t=x^{0}$, $r=x^{2}$, and $\theta=x^{3}$, while remaining independent of the azimuthal direction $\varphi=x^{1}$, reflecting the axisymmetric character of the perturbed geometry. This coordinate and notation scheme follows the framework adopted in Ref. chen2019gravitational . In the limit where the background spacetime remains static and spherically symmetric, the functions $\sigma$, $q_{2}$, and $q_{3}$ identically vanish. Therefore, in a linear perturbative analysis, these terms are introduced only at first order.
The next step involves reformulating the problem using the tetrad approach, where an orthonormal frame adapted to the spacetime described by metric (80) is introduced. This formalism streamlines the perturbative analysis by expressing tensorial quantities in terms of locally flat basis vectors. For further details on the implementation of this method, see Ref. chen2019gravitational
$e^{\mu}_{0}$ $=\left(e^{-\nu},\sigma e^{-\nu},0,0\right)\,,$
$e^{\mu}_{1}$ $=\left(0,e^{-\psi},0,0\right)\,,$
$e^{\mu}_{2}$ $=\left(0,q_{2}e^{-\mu_{2}},e^{-\mu_{2}},0\right)\,,$
$e^{\mu}_{3}$ $=\left(0,q_{3}e^{-\mu_{3}},0,e^{-\mu_{3}}\right)\,.$ (81)
Within this approach, indices corresponding to the tetrad basis are written in parentheses to clearly separate them from those tied to the coordinate system. The method involves translating all physical and geometrical quantities—originally formulated in terms of the metric $g_{\mu\nu}$—into a local inertial frame defined by the metric $\eta_{ab}$, through the application of a tetrad basis. Conventionally, $\eta_{ab}$ is identified with the Minkowski metric to simplify calculations. This redefinition allows vectors and tensors to be described using their components relative to the orthonormal frame rather than the coordinate basis
$A_{\mu}$ $=e_{\mu}^{a}A_{a}\,,\quad A_{a}=e_{a}^{\mu}A_{\mu}\,,$
$B_{\mu\nu}$ $=e_{\mu}^{a}e_{\nu}^{b}B_{ab}\,,\quad B_{ab}=e_{a}^{\mu}e_{b}^{\nu}B_{\mu\nu}\,.$ (82)
In the context of the tetrad formalism, the perturbed stress–energy tensor describing an anisotropic fluid is written in terms of components projected onto the orthonormal frame, yielding the following structure:
$\delta T_{ab}=$ $\,(\rho+p_{2})\delta(u_{a}u_{b})+(\delta\rho+\delta p_{2})u_{a}u_{b}$
$+(p_{1}-p_{2})\delta(x_{a}x_{b})+(\delta p_{1}-\delta p_{2})x_{a}x_{b}$
$+\delta p_{2}\eta_{ab}.$ (83)
By imposing the normalization and orthogonality conditions on the vectors $u^{\mu}$ and $x^{\mu}$—specifically, Eq. (75) together with $u^{\mu}x_{\mu}=0$—one obtains that all axial components of the perturbed stress–energy tensor vanish when expressed in the tetrad frame
$\delta T_{10}$ $=\delta T_{12}=\delta T_{13}=0\,.$ (84)
In the tetrad formalism, the Einstein field equations are recast in terms of tetrad components, leading to the following representation
$R_{ab}-\frac{1}{2}\eta_{ab}R=8\pi T_{ab}\,.$ (85)
Since the axial components of the perturbed stress–energy tensor vanish, the equation governing axial perturbations arises directly from the requirement $R_{ab}|_{\text{axial}}=0$. Through a sequence of algebraic manipulations—outlined in the Appendix of Ref. chen2019gravitational —one arrives at the corresponding master equation. The resulting effective potential for the axial gravitational perturbations is given by baruah2025quasinormal
$\mathcal{V}^{\,\text{t}}_{\text{metric}}(r)=g_{tt}(r)\left[\dfrac{2}{r^{2}}\left(\dfrac{1}{g_{rr}(r)}-1\right)+\dfrac{l(l+1)}{r^{2}}-\dfrac{1}{r\sqrt{g_{tt}(r)g_{rr}(r)}}\left(\dfrac{\mathrm{d}}{\mathrm{d}r}\sqrt{g_{tt}(r)g_{rr}^{-1}(r)}\right)\right],$ (86)
or, more explicitly
$\mathcal{V}^{\,\text{t}}_{\text{metric}}(r)=\left(1-\frac{2M}{r}\right)\left(\frac{l(l+1)}{r^{2}}-\frac{2(\ell r+3M)}{(\ell+1)r^{3}}\right).$ (87)
It is worth noting that in the limit $\ell\to 0$, the effective potential for odd–parity tensor perturbations recovers the familiar form associated with the Schwarzschild solution. With this groundwork in place, the study of greybody factors for tensorial perturbations can now be carried out as follows
$T^{\,\text{t}}_{b_{\text{metric}}}=\text{sech}^{2}\left[\frac{l(l+1)(\ell+1)-2\ell-1}{(2\omega)\left(4(\ell+1)^{3/2}M\right)}\right].$ (88)
Fig. 8 displays the greybody factors associated with tensor perturbations, $T^{\,\text{t}}_{b_{\text{metric}}}$, under two distinct conditions: the top panel illustrates the behavior as $\ell$ varies with $l$ fixed at 1, while the bottom panel shows results for different values of $l$ with $\ell$ held constant at 0.1. In each case, the outcomes are compared against the corresponding Schwarzschild baseline.
Figure 7: The effective potential $\mathcal{V}^{\,\text{t}}_{{\text{metric}}}$ is shown for different values of $\ell$ for $l=1$. Also, the Schwarzschild case is compared in this analysis.
Figure 8: The greybody factors $T^{\,\text{t}}_{b_{\text{metric}}}$ is displayed for different values of $\ell$ when keeping $l=1$ (the the top panel) and for different values of $l$ for a fixed value of $\ell=0.1$. For both cases, the Schwarzschild case is compared.
@@ IV.3.2 Vector perturbations
IV.3.3 Tensor perturbations
Following the analogy with the scalar and vector perturbations in the metric–affine framework, the effective potential is expressed as
$\begin{split}&\mathcal{V}^{\,\text{t}}_{{\text{met--aff}}}=\left(1-\frac{2M}{r}\right)\left(\frac{l(l+1)}{r^{2}}+\frac{6M(X-4)^{2}-2r(X-4)^{2}+8r\sqrt{-((X-4)(3X+4))}}{r^{3}(X-4)(3X+4)}\right).\end{split}$ (130)
Naturally, if $X\to 0$, we recover the effective potential for the Schwarzschild case odd parity (for the axial perturbations), namely, $\mathcal{V}^{\text{t}}_{\text{Schw}}=\left(1-\frac{2M}{r}\right)\left(\frac{l(l+1)}{r^{2}}-\frac{6M}{r^{3}}\right)$. Therefore, the greybody factors can be presented below
$\begin{split}&T^{\,\text{t}}_{b_{\text{met--aff}}}\geq{\mathop{\rm sech}\nolimits}^{2}\left[\int_{-\infty}^{+\infty}\frac{\mathcal{V}^{\,\text{t}}_{{\text{met--aff}}}}{2\omega}\mathrm{d}r^{*}\right]\\ &={\mathop{\rm sech}\nolimits}^{2}\left[\frac{(1-3l(l+1))X^{2}+8l(l+1)X+16l(l+1)-8\left(X+\sqrt{-((X-4)(3X+4))}-2\right)}{16M(3X+4)}\right].\end{split}$ (131)
Fig. 23 depicts the effective potential for tensor perturbations in the metric–affine formalism. It is observed that increasing $X$ results in a decrease in $\mathcal{V}^{\,\text{t}}_{{\text{met--aff}}}$. Meanwhile, Fig. 24 presents the greybody factors for varying values of $X$ with $l=1$ (top panel) and different values of $l$ for a fixed $X=0.1$, with both cases compared against the Schwarzschild scenario. Differently from the scalar and the vector perturbation, the increase of $X$ (for a fixed $l=1$) leads to an increase of greybody factors for the tensor perturbations.
Finally, in Fig. 25, the greybody factors for all bosonic perturbations are compared, revealing the following hierarchy: $T_{b_{\text{metric}}}^{\text{t}}>T_{b_{\text{met--aff}}}^{\text{t}}>\text{Schw}^{(\text{t})}>\text{Schw}^{(\text{v})}>T_{b_{\text{met--aff}}}^{\text{v}}>T_{b_{\text{metric}}}^{\text{v}}>\text{Schw}^{(\text{s})}>T_{b_{\text{met--aff}}}^{\text{s}}>T_{b_{\text{metric}}}^{\text{s}}$.
Additionally, for tensor perturbations, both $T_{b_{\text{metric}}}^{\text{t}}$ and $T_{b_{\text{met--aff}}}^{\text{t}}$ exhibit greater intensity than the corresponding Schwarzschild result—once again contrasting with the behavior observed for the other bosonic modes discussed in this work.
Figure 23: The effective potential $\mathcal{V}^{\,\text{t}}_{{\text{met--aff}}}$ is shown for different values of $X$ for $l=1$. Also, the Schwarzschild case is compared in this analysis.
Figure 24: The greybody factors $T^{\,\text{t}}_{b_{\text{met--aff}}}$ is displayed for different values of $X$ when keeping $l=1$ (the the top panel) and for different values of $l$ for a fixed value of $X=0.1$. For both cases, the Schwarzschild case is compared.
Figure 25: The comparison of the greybody factors for the bosonic case when $X=\ell=0.1$. The Schwarzschild case (for scalar, vector and tensor perturbations) is also present for the sake of comparison.